# Harpa: High-Rate Phase Association with Travel Time Neural Fields

### v2

@@ Results
In the known wave speed experiment, we used the SEG/EAGE 3-D overthrust model (?) (left plot in Fig. 1), rescaled to our experimental volume. It is remarkable that despite the complexity of the wave speed Harpa achieves near 100% association accuracy, even in the most challenging cases with many overlapping arrivals (green window in Fig. 4) where the ordinal number of arrivals is completely different from station to station. Another experiment with a known wave speed experiment is shown in Fig. S1.
Figure 8: Evaluation of Harpa with varying number of stations $(N)$ and sources $(M)$. Left to right, the four panels illustrate the impact on (a) association accuracy, (b) relative wave speed error, (c) relative earthquake location error, and (d) relative earthquake occurrence time error. Bars show the mean of 10 independent runs; error bars indicate standard deviation. For efficiency we used a smaller and simpler dataset, with up to $2$ random fields compared to $3$ in other experiments and the latent dimension $L=4$ instead of $L=6$ (See Materials and method for the data generation protocol). The mean $\mathsf{CF}$ is $0.80$.
We constructed a dataset of synthetic wave speeds by superimposing realizations of a Gaussian random field onto a vertical gradient wave speed (details in Materials and method). This is a challenging setting for association which exhibits strong spatial variation and a large five-fold dynamic range of wave speeds, typically between 5 km/s and 25 km/s. The test wave speed in this experiment comes from the same distribution (it is not included in the training set). In Materials and method we also report results for a different distribution of wave speeds obtained as random Gaussian perturbations.
@@
Distorted and out-of-distribution wave speeds
@@
Spurious and missing picks
@@
Discussion
Existing association algorithms operate under the assumption that the wave speed is approximately known and are applied in low-rate settings where the order of arrivals of waves from the different events coincides at all stations (?). A major challenge addressed by these methods is to sieve out spurious and missing arrivals which hamper association.
We explore a different regime where the arrivals are extremely dense in occurrence time, the ordinal number of arrivals at a station is a priori meaningless, and the wave speed is unknown. This regime is of great interest for earthquake monitoring but inaccessible to existing methods; see, for example, the density and mixing-up of arrivals that we successfully associate in Fig. 8, notwithstanding the unknown wave speed. The fact that we can still obtain accurate associations in situations of this complexity opens up new domains of application. Quoting from (?), “a comprehensive study of microseismicity can provide a valuable description of the geological medium properties and earthquake related processes in the investigated crustal volumes, such as, for instance, the identification and geometrical characterization of active fault structures.”.
Indeed, taming microseismicity would allow us to monitor the changes in time of the elastic material properties. One case in point is monitoring sustainable energy production facilities (?); another is understanding fault activity in earthquakes caused by “unconventional hydrocarbon development” (fracking); as Park and coauthors write (?), “Microearthquake activity can illuminate otherwise unknown faults, particularly when events are detected down to low magnitudes and are located precisely.”
From a high-level perspective, there are two reasons that make our proof of concept, Harpa, work. The first is that we take an inverse problems perspective and approximately recover continuous unknowns by interpreting the observed data as a probability measure. Indeed, in situations of such complexity, it is unlikely that statistical filtering methods can work well without explicitly using some model of wave propagation. Since we still only obtain (coarse) approximations of the relevant parameters, we cannot claim to solve the full travel time tomography problem, but it is conceivable that the resulting association and the approximate locations and wave speeds could be used as a warm start for a linearized or iterative method (?).
@@ Discussion
Materials and methods
Wave speed autoencoders
We discretize the wave speed into a 32 $\times$ 32 $\times$ 32 cube, i.e., $c\in\mathbb{R}^{32\times 32\times 32}$, and use these cubic images to train a standard CNN autoencoder in pytorch. Due to the low-dimensional structure of the dataset, the autoencoder exhibits strong generalization when an appropriate dimension of the latent code $z\in\mathbb{R}^{L}$ is selected. A large $L$ yields an autoencoder which accurately represents the wave speeds but also leads to slow SGLD inference. In our experiments we tested values of $L$ between 3 and 6. The generalization performance of the autoencoder is illustrated in Fig. S4.
@@
Training the travel time neural field
@@
We create a $(L+3)$-dimensional feature vector for each grid point by concatenating its coordinate and the corresponding wave speed latent code, which are all scaled in $[0,1]$ as the input to the neural field. We use a SIREN implicit network, a multilayer perceptron (MLP) with periodic activation functions (?). The trained SIREN model not only provides accurate predictions of the travel time at the grid points but also immediately yields an interpolation at arbitrary continuous coordinates. The scaling constant $w_{0}$ in the sinusoidal activation function is important and depends on the grid scaling; we used $w_{0}=15$.
Optimization
@@
We use the Python Optimal Transport (POT) package to compute the Wasserstein-2 distance, with $p=2$. We adopt the RMSProp-preconditioned version of the SGLD (?). Both the learning rate $(lr)$ and the noise factor $(\epsilon)$ are selected from $\{10^{-2},10^{-3},5\times 10^{-3},10^{-4}\}$. The convergence speed depends on the choice of hyperparameters. To manage the exploration-exploitation trade-off, we reduce the learning rate $(lr)$ to one-tenth of its original value once the loss decreases below a threshold $l^{*}$ which is selected from $\{10^{-2},10^{-3},10^{-4},10^{-6}\}$. To speed up training, we implement an epoch threshold. If the loss fails to decrease below $l^{*}$ before reaching this threshold, we reset all the parameters and restart the training process. We set the maximum number of trials to 10. When $\tau_{\max}$ is very large or $\mathsf{CF}$ is close to 00, pre-training $\tau$ helps achieve faster convergence.
Comparison between SGD and SGLD
@@
Here we show that due to non-convexity naive (stochastic) gradient descent converges to poor local minima, and that the issue is mitigated by SGLD. We compare SGD and SGLD on the fixed wave speed datasets described in Experiments. To visualize the optimization trajectory, we set the learning rate to $0.005$ and $\epsilon=0.0005$ (cf. (12)) for all epochs. Setting $\epsilon=0$ reduces SGLD to SGD; the corresponding performance is reported in the second row in Fig. S1. A visual comparison of the trajectories of the source estimates during optimization (Fig. S1, right) clearly shows that unlike SGD, SGLD easily escapes bad local minima.
Gaussian random field dataset
@@
In Figures 6, 7, and 8, the wave speeds are samples from the Gaussian random field dataset. To generate this low-dimensional wave speed dataset, we first created 20 three-dimensional Gaussian random fields as basis patterns in which the correlation is determined by the scale-invariant power spectrum $P(k)\sim k^{-\alpha/2}$ (?, ?). We use $\alpha=2.5$ in our experiments and rescale the pattern amplitudes to lie between 00 km/s and $15$ km/s; representative samples are shown in Fig. S3. For each sample, we randomly select up to 3 patterns and then add the selected patterns to a background wave speed (Fig. S2) with weights drawn uniform at random in $[-1,1]$. All wave speeds are clipped between $5$ km/s and $25$ km/s. In Fig. S4 we show that this dataset is effectively modeled by a 3D convolutional autoencoder.
Gaussian random perturbation dataset
@@
In Fig. S6 we report results for a different dataset in which we add between one and three random Gaussian perturbations to the background wave speed. These perturbations are defined as
@@
where $A$ is drawn uniformly at random from $[-25,25]$, $x_{0},y_{0},z_{0}$ uniformly at random from $[0,100]$ and $\sigma_{x},\sigma_{y},\sigma_{z}$ uniformly at random from $[10,50]$; samples are shown Fig. S5.
A Wasserstein misfit for unbalanced arrivals matching
@@
with $w_{p}$ defined in Equation (9). This mismatch can be optimized by sparsity-constrained optimal transport (?). In fact, in the 1D case it coincides with the loss of the unbalanced linear assignment problem
@@
Acknowledgments
Funding: C.S. and I.D. are supported by were supported by the European Research Council (ERC) Starting Grant 852821—SWING. M.V.d.H. is supported by the Simons Foundation under the MATH+X program, and Department of Energy, grant DE-SC0020345. Author contributions: C.S., M.V.d.H. and I.D. designed research, analyzed data, and wrote the paper. C.S. performed experiments. Competing interests: The authors declare that they have no competing interests. Data and materials availability: All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials. Code at https://github.com/DaDaCheng/phase_association.
Supplementary materials
Figs. S1 to S6
Supplementary Materials
Figure S1: Association results for $8$ sources, $20$ stations and $\mathsf{CF}=0.85$ with SEG/EAGE 3-D overthrust model as the known wave speed. Top row (SGLD): $\mathsf{ACC}=92.5\%$, $e_{l}=0.02s$. Bottom row (SGD): $\mathsf{ACC}=37.5\%$, $e_{l}=0.59s$. Symbols and encoding are as in Fig. 5; the right plot illustrates the source trajectories in the first 30 optimization epochs.
Figure S2: Background wave speed
Figure S3: Four samples of the 3D Gaussian random fields with $\alpha=2.5$.
Figure S4: A comparison between two test wave speeds unseen at training time and the corresponding outputs from an auto-encoder with $L=6$.
Figure S5: Four samples of wave speeds from the Gaussian perturbation dataset.
Figure S6: Association results for $8$ sources, $20$ stations and $\mathsf{CF}=0.79$ on Gaussian perturbation dataset. Harpa achieves $\mathsf{ACC}=96.3\%$ and $e_{l}=0.02$ s. Symbols and encoding are as in Fig. 6

### v3

@@ 3.2 Association via joint recovery
In the known wave speed experiment, we used the SEG/EAGE 3-D overthrust model [43] (left plot in Fig. 1), rescaled to our experimental volume. It is remarkable that despite the complexity of the wave speed Harpa achieves near 100% association accuracy, even in the most challenging cases with many overlapping arrivals (green window in Fig. 4) where the ordinal number of arrivals is completely different from station to station. Another experiment with a known wave speed experiment is shown in Supplementary information.
Figure 8: Evaluation of Harpa with varying number of stations $(N)$ and sources $(M)$. Left to right, the four panels illustrate the impact on (a) association accuracy, (b) relative wave speed error, (c) relative earthquake location error, and (d) relative earthquake occurrence time error. Bars show the mean of 10 independent runs; error bars indicate standard deviation. For efficiency we used a smaller and simpler dataset, with up to $2$ random fields compared to $3$ in other experiments and the latent dimension $L=4$ instead of $L=6$ (See Section 2 for the data generation protocol). The mean $\mathsf{CF}$ is $0.80$.
We constructed a dataset of synthetic wave speeds by superimposing realizations of a Gaussian random field onto a vertical gradient wave speed (details in Section 2). This is a challenging setting for association which exhibits strong spatial variation and a large five-fold dynamic range of wave speeds, typically between 5 km/s and 25 km/s. The test wave speed in this experiment comes from the same distribution (it is not included in the training set). In Section 2 we also report results for a different distribution of wave speeds obtained as random Gaussian perturbations.
@@
3.3 Distorted and out-of-distribution wave speeds
@@
3.4 Spurious and missing picks
@@
4 Discussion
Existing association algorithms operate under the assumption that the wave speed is approximately known and are applied in low-rate settings where the order of arrivals of waves from the different events coincides at all stations [5]. A major challenge addressed by these methods is to sieve out spurious and missing arrivals which hamper association.
We explore a different regime where the arrivals are extremely dense in occurrence time, the ordinal number of arrivals at a station is a priori meaningless, and the wave speed is unknown. This regime is of great interest for earthquake monitoring but inaccessible to existing methods; see, for example, the density and mixing-up of arrivals that we successfully associate in Fig. 8, notwithstanding the unknown wave speed. The fact that we can still obtain accurate associations in situations of this complexity opens up new domains of application. Quoting from [46], “a comprehensive study of microseismicity can provide a valuable description of the geological medium properties and earthquake related processes in the investigated crustal volumes, such as, for instance, the identification and geometrical characterization of active fault structures.”.
Indeed, taming microseismicity would allow us to monitor the changes in time of the elastic material properties. One case in point is monitoring sustainable energy production facilities [47]; another is understanding fault activity in earthquakes caused by “unconventional hydrocarbon development” (fracking); as Park and coauthors write [48], “Microearthquake activity can illuminate otherwise unknown faults, particularly when events are detected down to low magnitudes and are located precisely.”
From a high-level perspective, there are two reasons that make our proof of concept, Harpa, work. The first is that we take an inverse problems perspective and approximately recover continuous unknowns by interpreting the observed data as a probability measure. Indeed, in situations of such complexity, it is unlikely that statistical filtering methods can work well without explicitly using some model of wave propagation. Since we still only obtain (coarse) approximations of the relevant parameters, we cannot claim to solve the full travel time tomography problem, but it is conceivable that the resulting association and the approximate locations and wave speeds could be used as a warm start for a linearized or iterative method [49].
@@ 4 Discussion
Acknowledgments
C.S. and I.D. are supported by were supported by the European Research Council (ERC) Starting Grant 852821—SWING. M.V.d.H. is supported by the Simons Foundation under the MATH+X program, and Department of Energy, grant DE-SC0020345.
Author contributions
C.S, M.V.d.H. and I.D. design new models and experiments, analysed the data and wrote the paper. C.S performed the experiments.
Supplementary information
Figure 10: Association results for $8$ sources, $20$ stations and $\mathsf{CF}=0.85$ with SEG/EAGE 3-D overthrust model as the known wave speed. Top row (SGLD): $\mathsf{ACC}=92.5\%$, $e_{l}=0.02s$. Bottom row (SGD): $\mathsf{ACC}=37.5\%$, $e_{l}=0.59s$. Symbols and encoding are as in Fig. 5; the right plot illustrates the source trajectories in the first 30 optimization epochs.
Figure 11: Background wave speed
Figure 12: Four samples of the 3D Gaussian random fields with $\alpha=2.5$.
Figure 13: A comparison between two test wave speeds unseen at training time and the corresponding outputs from an auto-encoder with $L=6$.
Figure 14: Four samples of wave speeds from the Gaussian perturbation dataset.
Figure 15: Association results for $8$ sources, $20$ stations and $\mathsf{CF}=0.79$ on Gaussian perturbation dataset. Harpa achieves $\mathsf{ACC}=96.3\%$ and $e_{l}=0.02$ s. Symbols and encoding are as in Fig. 6
5 Model and training details
5.1 Wave speed autoencoders
We discretize the wave speed into a 32 $\times$ 32 $\times$ 32 cube, i.e., $c\in\mathbb{R}^{32\times 32\times 32}$, and use these cubic images to train a standard CNN autoencoder in pytorch. Due to the low-dimensional structure of the dataset, the autoencoder exhibits strong generalization when an appropriate dimension of the latent code $z\in\mathbb{R}^{L}$ is selected. A large $L$ yields an autoencoder which accurately represents the wave speeds but also leads to slow SGLD inference. In our experiments we tested values of $L$ between 3 and 6. The generalization performance of the autoencoder is illustrated in Fig. 13.
@@ 5.1 Wave speed autoencoders
5.2 Training the travel time neural field
@@
We create a $(L+3)$-dimensional feature vector for each grid point by concatenating its coordinate and the corresponding wave speed latent code, which are all scaled in $[0,1]$ as the input to the neural field. We use a SIREN implicit network, a multilayer perceptron (MLP) with periodic activation functions [8]. The trained SIREN model not only provides accurate predictions of the travel time at the grid points but also immediately yields an interpolation at arbitrary continuous coordinates. The scaling constant $w_{0}$ in the sinusoidal activation function is important and depends on the grid scaling; we used $w_{0}=15$.
5.3 Optimization
@@
We use the Python Optimal Transport (POT) package to compute the Wasserstein-2 distance, with $p=2$. We adopt the RMSProp-preconditioned version of the SGLD [50]. Both the learning rate $(lr)$ and the noise factor $(\epsilon)$ are selected from $\{10^{-2},10^{-3},5\times 10^{-3},10^{-4}\}$. The convergence speed depends on the choice of hyperparameters. To manage the exploration-exploitation trade-off, we reduce the learning rate $(lr)$ to one-tenth of its original value once the loss decreases below a threshold $l^{*}$ which is selected from $\{10^{-2},10^{-3},10^{-4},10^{-6}\}$. To speed up training, we implement an epoch threshold. If the loss fails to decrease below $l^{*}$ before reaching this threshold, we reset all the parameters and restart the training process. We set the maximum number of trials to 10. When $\tau_{\max}$ is very large or $\mathsf{CF}$ is close to 00, pre-training $\tau$ helps achieve faster convergence.
5.4 Comparison between SGD and SGLD
@@
Here we show that due to non-convexity naive (stochastic) gradient descent converges to poor local minima, and that the issue is mitigated by SGLD. We compare SGD and SGLD on the fixed wave speed datasets described in Section 3. To visualize the optimization trajectory, we set the learning rate to $0.005$ and $\epsilon=0.0005$ (cf. (12)) for all epochs. Setting $\epsilon=0$ reduces SGLD to SGD; the corresponding performance is reported in the second row in Fig. 10. A visual comparison of the trajectories of the source estimates during optimization (Fig. 10, right) clearly shows that unlike SGD, SGLD easily escapes bad local minima.
5.5 Gaussian random field dataset
@@
In Figures 6, 7, and 8, the wave speeds are samples from the Gaussian random field dataset. To generate this low-dimensional wave speed dataset, we first created 20 three-dimensional Gaussian random fields as basis patterns in which the correlation is determined by the scale-invariant power spectrum $P(k)\sim k^{-\alpha/2}$ [51, 52]. We use $\alpha=2.5$ in our experiments and rescale the pattern amplitudes to lie between 00 km/s and $15$ km/s; representative samples are shown in Fig. 12. For each sample, we randomly select up to 3 patterns and then add the selected patterns to a background wave speed (Fig. 11) with weights drawn uniform at random in $[-1,1]$. All wave speeds are clipped between $5$ km/s and $25$ km/s. In Fig. 13 we show that this dataset is effectively modeled by a 3D convolutional autoencoder.
5.6 Gaussian random perturbation dataset
@@
In Fig. 15 we report results for a different dataset in which we add between one and three random Gaussian perturbations to the background wave speed. These perturbations are defined as
@@
where $A$ is drawn uniformly at random from $[-25,25]$, $x_{0},y_{0},z_{0}$ uniformly at random from $[0,100]$ and $\sigma_{x},\sigma_{y},\sigma_{z}$ uniformly at random from $[10,50]$; samples are shown Fig. 14.
5.7 A Wasserstein misfit for unbalanced arrivals matching
@@
with $w_{p}$ defined in Equation (9). This mismatch can be optimized by sparsity-constrained optimal transport [53]. In fact, in the 1D case it coincides with the loss of the unbalanced linear assignment problem